OpenStudy (anonymous):

Solve cos2t+cost=0 for 0

OpenStudy (anonymous):

\[cos(2t) = cos^2t-sin^2t = 1- 2sin^2t\] \[cos(t)+cos^2t-sin^2t = 0\] \[cos(t)+1- 2sin^2t = 0\] \[cos(t) = 2sin^2t - 1\] \[cos(t) = -1(1-2sin^2t)\] I'm just playing around with the double-angle formula hoping to see something. I don't see it, but maybe you will... Sorry, I can't help with this much more...

OpenStudy (anonymous):

Here is the answer though. I am not sure how they arrived at it.

OpenStudy (anonymous):

Is that cos(2t) or \[cos^2t\] ? That will make a big difference; just checking...

OpenStudy (anonymous):

its is cos 2t...not cos^2 t

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

Okay. Well, since I'm not sure exactly how to solve this, if I were presented this on an exam I would draw my unit circle and plug in values for t that would make cos(2t) + cos(t) = 0. Just a fallback plan if you don't figure this out (although I don't advocate relying on this long-term).

OpenStudy (anonymous):

is sint+cost=1 a legitimate pythagorean/trig identity?

OpenStudy (anonymous):

perhaps i could use that?

OpenStudy (anonymous):

sin^2(x) + cos^2(x) = 1

OpenStudy (anonymous):

but if you square root it, could it work?

OpenStudy (anonymous):

You'd be taking the square-root of the whole left side which is different than sinx + cosx

OpenStudy (anonymous):

This might help: It used it religiously back in trig and cal2. http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

(although I never really spent enough time memorizing the trig identities..)

OpenStudy (anonymous):

alright, going to bed. good luck. try tagging some moderators or level-99 users with @username, like @mathbrz

OpenStudy (anonymous):

you can find them listed on the main page if they're online

OpenStudy (anonymous):

okay thanks

OpenStudy (aravindg):

Looks like your qn is already answered?

OpenStudy (anonymous):

no i dont have an answer yet.

OpenStudy (anonymous):

we were working towrds an answer, trying to explore our options

OpenStudy (anonymous):

i used the pythagorean identity to get 90 degrees and 270 degrees but turns out that i used it incorrectly.

OpenStudy (anonymous):

alright well im going to bed. it is late and i havent done my homework and tomorrow is going to be hard. if anyone figures out how to do this please post answers and explanations or hints on how to solve. thank you.